OpenStudy (anonymous):

How many solutions are there for the equation? x^2 + 16 = 8x

OpenStudy (anonymous):

just one

OpenStudy (anonymous):

\[x^2-8x+16=0\] \[(x-4)^2=0\] \[x=4\]

OpenStudy (anonymous):

what do you do after you get x=4 though

OpenStudy (kunal):

two answers... 4,4 .

OpenStudy (unklerhaukus):

you draw it \[f(x)=x^2-8x+16=(x-4)^2\]